OpenStudy (butterflyhope):

Will give medal: If a1 = 2 and An = 3a n - 1 + 1, then what is a3? I thought it may be An = 3(2) 3 - 1 but I don't know ; -; Thanks!

OpenStudy (abdullah1995):

what grade question is this ?

OpenStudy (butterflyhope):

8th-9th

OpenStudy (butterflyhope):

but I suck at math so lol

OpenStudy (amistre64):

typo .. a1=2 etc

OpenStudy (amistre64):

a1 = 2 and An = 3a n - 1 + 1, then what is a3? its just about following the rule a1 = 2 a2 = 3(a1) + 1 a3 = 3(a2)+1 = 3[3(a1) + 1] + 1 thats better :)

OpenStudy (butterflyhope):

Thank you :)

OpenStudy (amistre64):

yep